These items are taken from the financial statements of Kingbird, Inc. at December 31, 2017 $124,844 14,868 3,776 13,971 97,232 72,216 920 6,254 3,068 70,800 47,199 53,808 11,210 110,448 22,090 4,248 17,346 Accounts receivable Prepaid insurance Equipment Land Insurance expense Interest expense Common stock Retained earnings (January 1, 2017) Accumulated depreciation-buildings Accounts payable Notes payable Accumulated depreciation-equipment Interest payable Service revenue Prepare a classified balance sheet. Assume that $16,048 of the note payable will be paid in 2018. (List current assets in order of liquidity and property plant and equipment in order of land, buildings and

Kingbird, Inc. Balance Sheet Assets

Liabilities and Stockholders Equity
Answer #1

Kingbird, inc.

Balance sheet

December 31, 2017

Assets

Current Assets

Cash

13971

Accounts receivable

14868

Prepaid insurance

3776

Total current assets

32615

Plant, Property, and Equipment

Land

72216

Buildings

124844

Less: Accumulated Depreciation-Building

(53808)

71036

Equipment

97232

Less: Accumulated Depreciation -Equipment

(22090)

75142

218394

Total Assets

251009

Liabilities & Shareholder’s Equity

Current Liabilities

Accounts Payable

11210

Interest Payable

4248

Note Payable (current portion of long term debt)

16048

Total current liabilities

31506

Long Term Liabilities

Note Payable

94400

Total Liabilities

125906

Shareholders’ Equity

Common Stock

70800

Retained Earnings

54303

Total Shareholders’ Equity

125103

Total Liabilities & Shareholders’ Equity

251009

Long term liabilities =110448-16048 = 94400

Retained earnings = beginning balance + net income = 47199 + (17346-920-6254-3068) = 54303